Scan @ 11 Oct 2021 05:34 PM
---------------------------------------------
3.2
Calculate the mass of NHACI that should be dissolved in 525.00 mL of 0.1531 M
NH3 (K = 1.8 x 105) solution to prepare a buffer with a pH of 10.00. Assume that
the change in volume on addition of NHaCI is negligible.
(8)
1
Expert's answer
2021-10-13T04:47:03-0400

Molar mass of NH3 = 17.031 g/mol

0.1531× 17.031= 2.61

525÷2.61= 201.15g

1.8×105= 201.15× x

= 180000/201.15

= 894.85g